OpenStudy (zzr0ck3r):

Real analysis: Suppose that \(E\subset \mathcal{M}\), show that for each \(\epsilon>0\), there is an open set, \(O\), with \(E\subset O\), and \(\lambda(O-E)<\epsilon\)

OpenStudy (zzr0ck3r):

\(\mathcal{M}\), is the sigma algebra of Lebegue measurable sets, and \(\lambda\) is the Lebegue measure.

OpenStudy (zzr0ck3r):

Lebesgue*

OpenStudy (fibonaccichick666):

kind of looks like Lipschitz

OpenStudy (zzr0ck3r):

Lippelletz what?

OpenStudy (fibonaccichick666):

What class is this? (It's a theorem from calculus having to do with a specific type of limit)

OpenStudy (fibonaccichick666):

http://en.wikipedia.org/wiki/Lipschitz_continuity

OpenStudy (zzr0ck3r):

called math 511 at my university

OpenStudy (fibonaccichick666):

hmm, interesting. I'm sorry can't help, we didn't cover that in my analysis courses

OpenStudy (zzr0ck3r):

but tnx for trying:)

OpenStudy (abb0t):

@wio

OpenStudy (anonymous):

Do you know anything about E or O other than that they are in M?

OpenStudy (anonymous):

What about you construct O by taking the bounds of E and then adding \(\epsilon/2\) to both sides or something like that?

OpenStudy (anonymous):

I mean, adding a length of \(\epsilon / 2\)

OpenStudy (zzr0ck3r):

yeah, it seems super trivial, I just dont know how to show it. We dont know anything, but we do know that since E is measurable, its measure is the same as its outer measure, which is defined to be the infimum of the sums of I_n where I_n is an open cover of E, so right there I have an open set, that covers E. I just dont know how to talk about it.

OpenStudy (anonymous):

Unless \(E\) spans from infinity to infinity, it will have a max or minimum. Suppose it has a minimum \(a\). Then let \(O =(a-\epsilon,a-\epsilon/2)\cup E\). Isn't this an open set?

OpenStudy (anonymous):

By definition; \[ \lambda(E)= \inf\{ \sum_{n=1}^\infty Length(I_n)\quad |\quad E \subset \cup I_n \} \]

OpenStudy (anonymous):

Notice that we can take each \( I_n\) to be an open interval, because each interval [a,b] and each \( \epsilon >0\) we can find an open interval J such that \[ [a,b]\subset J=] a-\frac \epsilon 2 , b+ \frac \epsilon 2\\ Length[J]= Length[a,b] +\epsilon \] Also notice that \[ Length[I]= \lambda(I) \]

OpenStudy (anonymous):

By definition of inf, for every \(\epsilon >0\), there is a countable union of open intervals \(I_n\) that cover E and \[ \sum\lambda (I_n) \leq \lambda(E) + \epsilon\\ \]

OpenStudy (anonymous):

But \[ \lambda(\cup I_n) \leq \sum\lambda (I_n) \leq \lambda(E) + \epsilon\\ \]

OpenStudy (anonymous):

Let \[ O= \cup I_n \] then \[ E\subset O\\ \lambda(O) \leq \lambda(E) +\epsilon\\ \lambda(O\setminus E)\leq \epsilon \]

OpenStudy (zzr0ck3r):

where would the strict inequality come from? I'm thinking the definition of inf, which is what I was having a problem with. I never really knew how to talk about our cover explicitly and now I do. Ty again. can you come teach my class?

OpenStudy (anonymous):

You can actually take the inequality to be strict. If a is the inf of a bunch of numbers, then for every \( \epsilon > 0\) and if you consider \( a +\epsilon >a \), then there must be one number of these bunch of numbers that is strictly less than \( a +\epsilon \) Thanks for trusting me to teach your class. I love measure theory, I spent all my life teaching it and doing research about it. Do not hesitate to ask me for help in the future.

OpenStudy (anonymous):

Let\[ O= \cup I_n \] then \[ E\subset O\\ \lambda(O) < \lambda(E) +\epsilon\\ \lambda(O\setminus E)< \epsilon \]